OpenStudy (anonymous):

Solve 2x + 3y + z = 12 using matrices. 3x + y – 4z = -5 -4x – y + 3z = 6

OpenStudy (loser66):

Your idea?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

could you help me with this one too Solve 2x – 4y + 3z = -8 using matrices. x + 3y – 2z = 9 3x + 2y + z = 13

OpenStudy (loser66):

I don't do your homework, you do it. If you stuck at somewhere, I help. You MUST have something, if you don't, how can I help?

OpenStudy (anonymous):

I came down to these four choices I am not sure which one is correct: (2, 2, 2) (0, 0, 12) (-2, -2, -2) (-2, 5, 1)

OpenStudy (anonymous):

for the fist quesstion:

OpenStudy (anonymous):

is it a or b

OpenStudy (perl):

$$ \Large { 2x + 3y + z = 12 \\ 3x + y ~–~ 4z = -5 \\ -4x~– y + 3z = 6\\ \\~\\\iff \\~\\ \begin{bmatrix} 2 &3 & 1 \\ 3&1&-4\\ -4&-1&3 \end{bmatrix} \begin{bmatrix} x\\ y\\ z \end{bmatrix} = \begin{bmatrix} 12\\ -5\\ 6 \end{bmatrix} }$$

OpenStudy (anonymous):

?

OpenStudy (perl):

theres a few ways you can solve this, you can make an augmented matrix , or solve it by multiplying both sides by A inverse

OpenStudy (anonymous):

I go 4 choices that I think one has to be correct

OpenStudy (perl):

It's against open study policy to give out answer, but I can guide you

OpenStudy (anonymous):

ok